OpenStudy (anonymous):

Let \(A\) be an \(n\times n\) matrix with \(a_{i,j}=\min\{i,j\}\) and \(1\le i\le n,~1\le j\le n\). Find the determinant of \(A\).

OpenStudy (anonymous):

Note: This question was adapted from this year's Putnam exam, which asked for the determinant of \(A\) but with \(a_{i,j}=\dfrac{1}{\min\{i,j\}}\). Also a fun problem.

OpenStudy (anonymous):

If the notation isn't familiar, the matrix is \[A=\begin{pmatrix} 1&1&1&\cdots&1&1&1\\ 1&2&2&\cdots&2&2&2\\ 1&2&3&\cdots&3&3&3\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1&2&3&\cdots&n-2&n-2&n-2\\ 1&2&3&\cdots&n-2&n-1&n-1\\ 1&2&3&\cdots&n-2&n-1&n \end{pmatrix}\]

OpenStudy (anonymous):

lol was trying to write that

OpenStudy (anonymous):

the only thing i note that it sound symmetric to me :)

OpenStudy (anonymous):

I don't have the precise solution for this version of the question (I'd actually started solving the original one while mistaking \(\min\{\}\) for \(\dfrac{1}{\min\{\}}\)... -_-)

OpenStudy (anonymous):

... but I do have one method.

OpenStudy (anonymous):

My approach involved writing \(A\) as a triangular matrix \(\bar{A}\) using a set of somewhat involved row operations, then finding the product of the diagonals.

hartnn (hartnn):

isn't the determinant always 1 irrespective of order ?

hartnn (hartnn):

with your approach you'd get all 1s in the diagonal, so yeah :)

OpenStudy (anonymous):

i wish its a triangle matrices so det would be n! :P (just saying lol )

hartnn (hartnn):

wondering how the problem with \(a_{i,j}=\dfrac{1}{\min\{i,j\}}\) would work :O

ganeshie8 (ganeshie8):

\[\prod_{i=2}^n\dfrac{-1}{i(i-1)} \]

hartnn (hartnn):

-1/2, 1/12, -1/144,1/2880,... is that what we get by that product .. .

hartnn (hartnn):

yep, thats correct....how did u get that /

ganeshie8 (ganeshie8):

just used ur hint..

ganeshie8 (ganeshie8):

we do \(n-1\) row operations : \(R_i - R_{i-1}\) \(2\le i \le n\)

OpenStudy (anonymous):

cute :3

ganeshie8 (ganeshie8):

the inverse of this matrix looks interesting its both symmetric and tridiogonal http://www.wolframalpha.com/input/?i=inverse+%7B%7B1%2C1%2C1%2C1%7D%2C%7B1%2C2%2C2%2C2%7D%2C%7B1%2C2%2C3%2C3%7D%2C%7B1%2C2%2C3%2C4%7D%7D

ganeshie8 (ganeshie8):

i have tested upto n=5 i feel the pattern continues

OpenStudy (anonymous):

yeah it looking nice

OpenStudy (anonymous):

eh , i forgot all interesting things in matrices :P

hartnn (hartnn):

n-1 column operations are also needed right ? Rn - R(n-1) and then Cn - C(n-1)

hartnn (hartnn):

*to make it a diagonal matrix

ganeshie8 (ganeshie8):

Oh yes, we can try finding the inverse step by step to see how it is tridiogonal but for the determinant, triangular form is sufficient right ?

hartnn (hartnn):

yes, correct

OpenStudy (anonymous):

@ganeshie8 I'd ended up with the same, \[\frac{(-1)^{n+1}}{n!(n-1)!}\] Row operations and general procedure: Replace the \(j\)th column entry of the \(i\)th row, \(R_i\), with \(R_i-(i-1)!R_1\). This makes the last column all zero except for the first entry which is 1. A cofactor expansion along this column gives \[\det(A)=(-1)^{n+1}\left(\prod_{k=1}^nk!\right)^{-1}\det(\bar{A})\] where \(\bar{A}\) was used to denote the submatrix containing \(a_{i,j}\) with \(2\le i\le n\) and \(1\le j\le n-1\). All the while, \(\bar{A}\) is triangular and so \(\det(\bar{A})\) is the product of the diagonal entries, and that product looks like \[\det(\bar{A})=\prod_{k=1}^{n-1}\left(\frac{(k+1)!}{k}-k!\right)\] and so \[\begin{align*}\det(A)&=(-1)^{n+1}\prod_{k=1}^{n-1}\frac{\dfrac{(k+1)!}{k}-k!}{(k+1)!}\\\\ &=(-1)^{n+1}\prod_{k=1}^{n-1}\frac{\dfrac{k+1}{k}-1}{k+1}\\\\ &=(-1)^{n+1}\prod_{k=1}^{n-1}\frac{1}{k(k+1)}\\\\ &=\frac{(-1)^{n+1}}{(n-1)!n!} \end{align*}\] Good to know I got that one right!

ganeshie8 (ganeshie8):

Wow! looks nice :) I think reducing the matrix from the bottom right corner would make it easy though @SithsAndGiggles I am not able to find putnam 2014 problems online. how did you get this problem ?

OpenStudy (anonymous):

I had the luxury of taking it yesterday morning at my school. I can post the other problems if anyone is interested (unless there's some legal issue surrounding that?).

ganeshie8 (ganeshie8):

XD I wish you all the good luck for top place! let us know when they announce the results :) There should not be any issues as they will be publishing problems and solutions shortly http://math.scu.edu/putnam/announcecJan.html

OpenStudy (anonymous):

hhehe , never had something like this

OpenStudy (anonymous):

I remember reading somewhere that the annual mean score is 0 out of 120, so I'm not too hopeful... and I was the only one taking the test this year.

ganeshie8 (ganeshie8):

0 out of 120 ?! negative marks for wrong attempts is it! I read something like they rank based on the overall performance of a 3 member team ? @Marki looks artofproblemsolving is super fast in releasing keys xD

hartnn (hartnn):

Patniss and Keeta in 7th Annual Putnal Games! ROFL

hartnn (hartnn):

*75th

OpenStudy (anonymous):

since its art works ;) so no doubt its fast